# Bayesian Optimisation with Unknown Hyperparameters: Regret Bounds Logarithmically Closer to Optimal

### v1

@@ 3 Length scale Balancing
The algorithm starts in line 1 by initialising the set of candidates to just the upper bound $\theta_{0}$. Later on, in lines 14-16, new candidates are introduced using the candidate-suggesting function $q(\cdot)$ at a pace dictated by the growth function $g(t)$. Typically in aggregation schemes, each one of the base algorithms is run in isolation. However, there is nothing preventing us from making them share the data and as such, selecting a base algorithm in our case simply amounts to choosing the length scale value we will use to fit the GP model at a given time step $t$, which is done in line 3. This choice is done by the regret-balancing rule $\theta_{t}=\arg\min_{\theta\in\Theta_{t}}R^{\theta}(|S_{t}^{\theta}|+1)$, with $R^{\theta}(\cdot)$ defined as in Theorem 2.2 and $S_{t}^{\theta}$ being the set of iterations before $t$ at which length scale value $\theta$ was chosen. Note that this rule implies that lower length scale values will be selected less frequently than higher values, as their regret bounds grow faster. After that, in line 4, the algorithm utilises the acquisition rule dictated by a model fitted with the selected length scale value to find the point to query next, $\bm{x}_{t}$. The idea is that, occasionally, $\theta_{t}$ will be set to one of the smaller values from $\Theta_{t}$ and, if that results in finding significantly better function values, then the rejection mechanism in lines 9-12 will remove longer length scales from the set of considered values, $\Theta_{t}$. Otherwise, we will keep all of the length scales and try again after some number of iterations. We now proceed to derive a regret bound for our developed algorithm.
@@ 4 Regret Bound and Proof Sketch
Let us use confidence parameters of $\xi_{t}=2\sigma_{N}^{2}\log\left(\hat{d}\ln(g(t))\pi^{2}t^{2}\right)-\log 3\delta$ and $\beta^{\theta}_{t}=B(\theta,N^{\star})+\sigma_{N}\sqrt{2(\gamma^{\theta}_{t-1}+1+\ln(2/\delta))}$, then Algorithm 1, achieves with probability at least $1-\delta$, the cumulative regret $R_{T}$ of the algorithm admits the following bound:
$R_{T}=\mathcal{O}\left(\left(t_{0}+\iota\right)B^{\star}+\left(R^{\theta^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{\hat{d}}\sqrt{\iota}+\iota\right)\right),$
where $t_{0}=g^{-1}\left(e^{-1/\hat{d}}\theta_{0}/\theta^{\star}\right)$ and $\iota=\hat{d}\ln g(T)$. We denote $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ in case of a $\nu$-Matérn kernel.
@@
(sketch) We provide a sketch of the result here and defer the proof to Appendix C.
Let us denote by $t_{0}$ the iteration at which the first well-specified length scale ($\theta\leq\theta^{\star}$) is added to the candidate set in line 15. This will happen at the first iteration after $g^{-1}(\frac{\theta_{0}}{\theta^{\star}})$, where the condition in line 14 will trigger. Given the ratios between consecutive candidates suggested by $q(\cdot)$, we get that $t_{0}=\lceil g^{-1}(\frac{\theta_{0}}{\theta^{\star}}e^{-1/d})\rceil$, in case of an RBF kernel, and, $t_{0}=\lceil g^{-1}(\frac{\theta_{0}}{\theta^{\star}}e^{-1/(2\nu+d)})\rceil$, in case of a $\nu$-Matérn kernel. On iterations up to $t_{0}$, we can potentially suffer the highest as possible, thus the cumulative regret can be bounded as:
@@
$\tilde{R}_{T}=\sum_{t\in\mathcal{T}}r_{t}+\sum_{t\notin\mathcal{T}}r_{t}\leq 2B|\mathcal{T}|+\sum_{t\notin\mathcal{T}}r_{t}\leq 2B^{\star}q^{-1}(\frac{\theta_{0}}{g(T)})+\sum_{t\notin\mathcal{T}}r_{t},$
where the second inequality comes from the fact that we cannot reject more candidates than we have introduced in total. The remaining thing to do is to bound $\sum_{t\notin\mathcal{T}}r_{t}$. This expression is the cumulative regret of the iterations, where no candidates are rejected and where at least one of the well-specified candidates has been introduced. First, we show that, with a probability of at least $1-\delta$, the well-specified candidate introduced at $t_{0}$ will not be rejected. Second, since no other candidates are rejected at iterations $t\notin\mathcal{T}$, it means that the function values achieved at those iterations cannot be too different from the ones achieved when using $\hat{\theta}$. Using this fact, we arrive at a statement:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\left(R^{\hat{\theta}}(|S_{t}^{\hat{\theta}}|)+\sqrt{T\xi_{T}}\right)\left(\sqrt{\sum_{\theta\in\Theta_{T}}\frac{|S_{t}^{\theta}|}{|S_{t}^{\hat{\theta}}|}}+q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)\right).$
The rest of the proof consists of bounding $\frac{|S_{t}^{\theta}|}{|S_{t}^{\hat{\theta}}|}$, which can be done due to the selection rule in line 3.
Table 1: Comparison of sub-optimality for A-GP-UCB and LB-GP-UCB for fixed growth functions $g(\cdot)$ and $b(\cdot)$. We denote $\hat{d}=d$ in case of an RBF kernel and $\hat{d}=d+2\nu$ for a $\nu$-Matérn kernel.
Algorithm Optimality $R_{T}/R^{\theta^{\star}}(T)$
@@
A-GP-UCB [6] $\mathcal{O}(g(T)^{\hat{d}})$ $\mathcal{O}(b(T)g(T)^{\hat{d}})$
LB-GP-UCB / LNB-GP-UCB (ours) $\mathcal{O}({\hat{d}}\ln g(T))$ $\mathcal{O}({\hat{d}}\ln g(T)\ln b(T))$
t Optimality Observe that for a fixed choice of growth function, we get $R_{T}/R^{\theta^{*}}(T)=\mathcal{O}(\hat{d}\ln g(T))$. This is an improvement compared to A-GP-UCB achieving $R_{T}/R^{\theta^{*}}(T)=\mathcal{O}(g(T)^{\hat{d}})$. As such the bound of our algorithm is significantly closer to the optimal bound than the one of A-GP-UCB. The faster $g(\cdot)$ is increasing, the quicker we will be able to find the first well-specified candidate, which will decrease the term $t_{0}B^{\star}$ in the bound. At the same time, it will increase all the terms depending on $g(T)$, but as our bound only scales with $\hat{d}\log g(T)$, we are able to select much more aggressive growth functions than A-GP-UCB, whose bound scales with $g(T)^{\hat{d}}$. In the Experiments section we compare the performance of LB-GP-UCB and A-GP-UCB using different growth functions $g(t)$ and show that the former algorithm is much more robust to the choice of $g(t)$.
Extension to unknown $N^{\star}$ As we discussed before, LB-GP-UCB requires us to know the optimal RKHS norm $N^{\star}$. However, we can easily extend the algorithm to handle the case of unknown $N^{\star}$, which we do in Appendix D. In Algorithm 3 we present Length scale and Bound Balancing (LNB) — an algorithm, which in addition to having candidates for $\theta$ also maintains a number of candidates for $N^{\star}$. As such, it requires us to specify another growth function $b(t)$ for exploring new norm values as well as the initial RKHS norm $B_{0}$. We prove its cumulative regret bound in Theorem D.1. We can similarly derive the suboptimality gap for our algorithm in this case. We display it in Table 1 together with the gap of A-GP-UCB. We can see that in this setting, we also achieve an improvement.
@@ 5 Experiments
We now evaluate the performance of our algorithm on multiple synthetic and real-world functions. To run experiments we used the compute resources listed in Appendix E.1 and implemented based on the codebase of [2], which uses the BoTorch package [5, 30].We used the UCB acquisition function and we compared different techniques for selecting the length scale value. For all experiments, we used isotropic $\nu$-Matern kernel with $\nu=2.5$. We standardise the observations before fitting the GP model and as such keep the kernel outputscale fixed to $1.0$. The first baseline we compare against is MLE, where the length scale value is optimised using a multi-start L-BFGS-B method [24] (the default BoTorch optimiser [5]) after each timestep by maximising the marginal likelihood for the data collected so far. The next baseline is MCMC, where we employ a fully Bayesian treatment of the unknown length scale value using the NUTS sampler [18], which we implemented using Pyro [7]. We use BoTorch’s default hyperprior ($\theta\sim\Gamma(3,6)$) and to select a new point we optimise the expected acquisition function under the posterior samples as described by [13]. We also compare against A-GP-UCB. To achieve a fair comparison, we used the same growth function $g(t)=\max\{t_{0},\sqrt{t}\}$ for both LB-GP-UCB and A-GP-UCB across all experiments, where $t_{0}$ was selected so that at least 5 candidates are generated for $g(1)$. We study the impact of this choice in the ablation section. We used 10 initial points for each algorithm unless specified otherwise. To select upper bound $\theta_{0}$ for A-GP-UCB and LB-GP-UCB we fitted a length scale to initial data points with MLE (and we did not use MLE after that). We present the results in Figure 3 below. We now describe each benchmark problem we used more in detail.
@@
Berkenkamp Toy Problem We start with a toy problem proposed by the same paper that proposed the A-GP-UCB algorithm [6]. We showed a plot of this one-dimensional function in Figure 1. In the right side of the domain, the function appears is smoother than on the left side. In this problem, we only use three initial points, to benchmark the ability of algorithms to escape from the local optimum in the right side of the domain. MLE and MCMC can be easily misled towards too-long length scale values, which causes them to get stuck in the local optimum. Both A-GP-UCB and LB-GP-UCB quickly find the optimal solution, however, due to over-exploration, the cumulative regret of A-GP-UCB grows faster than that of LB-GP-UCB.
@@
Michalewicz Synthetic Function As a next benchmark, we evaluate our algorithm on the Michalewicz synthetic function, which has been designed to be challenging while using MLE for fitting hyperparameters, because it exhibits different degrees of smoothness throughout its domain. In the histogram in Figure 2, we compare the selected length scale value with an estimate of the optimal length scale $\hat{\theta}^{*}$. We produce this estimate by sampling ten thousand points uniformly through the domain and fitting a length scale value by maximum likelihood to the points with the top 1% of objective values. In this way, we are able to capture the length scale value that produces a good model for the Michalewicz function around the optimum, where it is least smooth. We can see LB-GP-UCB selects lengthscale value closer to $\hat{\theta}^{*}$ and as a result outperforms other baselines in terms of regret.
@@
Material Design Problems We utilise material design tasks proposed by [16] and [27]. At each time step, the algorithm can choose which material configuration to try, and observe the objective value, which corresponds to a given material optimisation criterion. As material design problems are known to exhibit a needle-in-a-haystack behaviour [31], on both benchmarks MLE and MCMC get stuck at a suboptimal solution and their best regret does not fall beyond a certain value. A-GP-UCB is able to quickly find low-regret solutions on the AGNP benchmark, but struggles on the Crossedbarrel problem and underperforms in terms of cumulative regret. On the contrary, LB-GP-UCB performs well across both benchmark problems and across both regret metrics.
@@
Ablation on $g(t)$ To test robustness of LB-GP-UCB, we evaluate it on Michalewicz function together with A-GP-UCB for different choices of $g(t)$. We try functions of form $g(t)=\max(t_{0},t^{a})$ for $a\in(0.25,0.5,0.75)$. In Figure 4 we plot the final performance of the algorithms after $N=250$ steps as well as the distribution of selected length scale values. We see that A-GP-UCB is very sensitive to the selection of growth function $g(t)$, whereas our algorithm selects similar length scale values regardless of $g(t)$, which results in consistently good best regret results. We can see that LB-GP-UCB typically selects values around $\hat{\theta}^{*}$ for different growth functions, whereas A-GP-UCB decreases its length scale beyond $\hat{\theta}^{*}$, resulting in slower convergence.
@@ 6 Related Work
We already mentioned the work proposing A-GP-UCB [6], which so far has been the only work providing the guarantees on BO with unknown hyperparameters, where only an upper bound on the optimal length scale value is known. The work of [34] addresses the problem, where, in addition, a lower bound on the optimal length scale is known, however, the regret bound of the algorithm they propose scales with the $\gamma(k^{\theta_{L}})$ of the smallest possible length scale $\theta_{L}$, making it no better than a naïve algorithm always selecting $\theta_{L}$. [8] studied the problem of solving BO, when the kernel function is misspecified, however, provided no method for finding the well-specified kernel function. [10] derived predictive guarantees for GP with unknown hyperparameters, provided that hyperpriors on those hyperparameters are known. However, the authors do not provide any BO algorithm based on their results. There have also been a number of works[13, 20, 25, 28] that tackled the problem of BO with unknown hyperparameters but did not provide a theoretical analysis of the used algorithm. One of the recent works [37], showed that choosing kernel decompositions (which can be treated as another hyperparameter) randomly can outperform learning them by MLE. Some of earlier works[14, 33] viewed BO with unknown hyperparameters as meta-learning or transfer learning problem, where a large dataset is available for pre-training. In our problem setting, we do not assume access to any such pre-training data. Within this work, we considered a frequentist problem setting, where the black-box function is arbitrarily selected from some RKHS. While there are no guarantees for MLE to be consistent in such a setting, if we were to assume a Bayesian setting and put a GP prior on the black box, statistical literature derived asymptotic consistency results [4, 21, 23, 26] for MLE of kernel hyperparameters, including length scale.

### v3

@@ 3 Length scale Balancing
The algorithm starts in line 1 by initialising the set of candidates to just the upper bound $\theta_{0}$. Later on, in lines 14-16, new candidates are introduced using the candidate-suggesting function $q(\cdot)$ at a pace dictated by the growth function $g(t)$. Typically in aggregation schemes, each one of the base algorithms is run in isolation. However, there is nothing preventing us from making them share the data and as such, selecting a base algorithm in our case simply amounts to choosing the length scale value we will use to fit the GP model at a given time step $t$, which is done in line 3. This choice is done by the regret-balancing rule $\theta_{t}=\arg\min_{\theta\in\Theta_{t}}R^{\theta}(|S_{t}^{\theta}|+1)$, with $R^{\theta}(\cdot)$ defined as in Theorem 2.3 and $S_{t}^{\theta}$ being the set of iterations before $t$ at which length scale value $\theta$ was chosen. Note that we only need to know the scaling of the bound up to a constant. This rule implies that lower length scale values will be selected less frequently than higher values, as their regret bounds grow faster. After that, in line 4, the algorithm utilises the acquisition rule dictated by a model fitted with the selected length scale value to find the point to query next, $\bm{x}_{t}$. The idea is that, occasionally, $\theta_{t}$ will be set to one of the smaller values from $\Theta_{t}$ and, if that results in finding significantly better function values, then the rejection mechanism in lines 9-12 will remove longer length scales from the set of considered values, $\Theta_{t}$. Otherwise, we will keep all of the length scales and try again after some number of iterations. We now proceed to derive a regret bound for our developed algorithm.
@@ 4 Regret Bound and Proof Sketch
Let us use confidence parameters of $\xi_{t}=2\sigma_{N}^{2}\log\left(d\ln(g(t))\pi^{2}t^{2}\right)-\log 3\delta$ and $\beta^{\theta}_{t}=B(\theta,N^{\star})+\sigma_{N}\sqrt{2(\gamma^{\theta}_{t-1}+1+\ln(2/\delta))}$, then with probability at least $1-\delta$, the cumulative regret $R_{T}$ of the Algorithm 1 admits the following bound:
$R_{T}=\mathcal{O}\left(\left(t_{0}+\iota\right)B^{\star}+\left(R^{\theta^{\star}}(T)+\sqrt{T\xi_{T}}\right)\left(\left(\frac{\theta_{0}}{\theta^{\star}}\right)^{d}d\ln\frac{\theta_{0}}{\theta^{\star}}+\iota\right)\right),$
where $t_{0}=g^{-1}\left(e^{-1/d}\theta_{0}/\theta^{\star}\right)$ and $\iota=d\ln g(T)$.
@@
(sketch) We provide a sketch of the result here and defer the proof to Appendix D.
Let us denote by $t_{0}$ the iteration at which the first well-specified length scale ($\hat{\theta}\leq\theta^{\star}$) is added to the candidate set in line 15. This will happen at the first iteration after $g^{-1}(\frac{\theta_{0}}{\theta^{\star}})$, where the condition in line 14 will trigger. Given the ratios between consecutive candidates suggested by $q(\cdot)$, we get that $t_{0}=\lceil g^{-1}(\frac{\theta_{0}}{\theta^{\star}}e^{-1/d})\rceil$. On iterations up to $t_{0}$, we can potentially suffer the highest as possible, thus the cumulative regret can be bounded as:
@@
$\tilde{R}_{T}=\sum_{t\in\mathcal{T}}r_{t}+\sum_{t\notin\mathcal{T}}r_{t}\leq 2B^{\star}|\mathcal{T}|+\sum_{t\notin\mathcal{T}}r_{t}\leq 2B^{\star}q^{-1}(\frac{\theta_{0}}{g(T)})+\sum_{t\notin\mathcal{T}}r_{t},$
where the second inequality comes from the fact that we cannot reject more candidates than we have introduced in total. The remaining thing to do is to bound $\sum_{t\notin\mathcal{T}}r_{t}$. This expression is the cumulative regret of the iterations, where no candidates are rejected and where at least one of the well-specified candidates has been introduced. We can bound this term using a similar strategy as in [33]. First, we show that, with a probability of at least $1-\delta$, the well-specified candidate introduced at $t_{0}$ will not be rejected. Second, since no other candidates are rejected at iterations $t\notin\mathcal{T}$, it means that the function values achieved at those iterations cannot be too different from the ones achieved when using $\hat{\theta}$. Using this fact, we arrive at a statement:
$\sum_{t\notin\mathcal{T}}r_{t}\leq\left(R^{\hat{\theta}}(|S_{t}^{\hat{\theta}}|)+\sqrt{T\xi_{T}}\right)\left(\sum_{\theta\in\mathcal{M}_{0}}\sqrt{\frac{|S_{t}^{\theta}|}{|S_{t}^{\hat{\theta}}|}}+q^{-1}\left(\frac{\theta_{0}}{g(T)}\right)\right),$
where $\mathcal{M}_{0}$ is the set of misspecified length scale values that were chosen at least once after $t_{0}$. The rest of the proof consists of bounding $\frac{|S_{t}^{\theta}|}{|S_{t}^{\hat{\theta}}|}$, which can be done due to the selection rule in line 3.
Optimality In Appendix H, we show that for a fixed choice of growth function, we get $R_{T}/R^{\theta^{\star}}(T)=\mathcal{O}(d\ln g(T))$. This is an improvement compared to A-GP-UCB achieving $R_{T}/R^{\theta^{\star}}(T)=\mathcal{O}(g(T)^{d})$. As such the bound of our algorithm is significantly closer to the optimal bound than the one of A-GP-UCB. The faster $g(\cdot)$ is increasing, the quicker we will be able to find the first well-specified candidate, which will decrease the term $t_{0}B^{\star}$ in the bound. At the same time, it will increase all the terms depending on $g(T)$, but as our bound only scales with $d\log g(T)$, we are able to select much more aggressive growth functions than A-GP-UCB, whose bound scales with $g(T)^{d}$. In the Experiments section we compare the performance of LB-GP-UCB and A-GP-UCB using different growth functions $g(t)$ and show that the former algorithm is much more robust to the choice of $g(t)$.
Table 1: Comparison of optimality for A-GP-UCB and LB-GP-UCB for fixed functions $g(\cdot)$ and $b(\cdot)$. $R^{\star}(T)$ refers to the scaling of the regret bound of an oracle optimiser, knowing the optimal hyperparameters. See Appendix H for more details.
Algorithm Optimality $R_{T}/R^{\star}(T)$
@@
A-GP-UCB [6] $\mathcal{O}(g(T)^{d})$ $\mathcal{O}(b(T)g(T)^{d})$
LB-GP-UCB / LNB-GP-UCB (ours) $\mathcal{O}({d}\ln g(T))$ $\mathcal{O}({d}\ln g(T)\ln b(T))$
Extension to unknown $N^{\star}$ As we discussed before, LB-GP-UCB requires us to know the initial RKHS norm $N^{\star}$. However, we can easily extend the algorithm to handle the case of unknown $N^{\star}$, which we do in Appendix F. In Algorithm 3 we present Length scale and Bound Balancing (LNB) — an algorithm, which in addition to having candidates for $\theta$ also maintains a number of candidates for $N^{\star}$. As such, it requires us to specify another growth function $b(t)$ for exploring new norm values as well as the initial RKHS norm $B_{0}$. We prove its cumulative regret bound in Theorem F.1. We can similarly derive the suboptimality gap for our algorithm in this case. We display it in Table 1 together with the gap of A-GP-UCB. We can see that in this setting, we also achieve an improvement.
@@ 5 Experiments
We now evaluate the performance of our algorithm on multiple synthetic and real-world functions. To run experiments we used the compute resources listed in Appendix I.1 and implemented based on the codebase of [2], which uses the BoTorch package [5, 34]. We open-source our code [^2] https://github.com/JuliuszZiomek/LB-GP-UCB. We used the UCB acquisition function and we compared different techniques for selecting the length scale value. For all experiments, we used isotropic $\nu$-Matérn kernel with $\nu=2.5$. We standardise the observations before fitting the GP model and as such keep the kernel outputscale fixed to $1.0$. The first baseline we compare against is MLE, where the length scale value is optimised using a multi-start L-BFGS-B method [25] (the default BoTorch optimiser [5]) after each timestep by maximising the marginal likelihood for the data collected so far. The next baseline is MCMC, where we employ a fully Bayesian treatment of the unknown length scale value using the NUTS sampler [19], which we implemented using Pyro [7]. We use BoTorch’s default hyperprior ($\theta\sim\Gamma(3,6)$) and to select a new point we optimise the expected acquisition function under the posterior samples as described by [13]. We also compare against A-GP-UCB. To achieve a fair comparison, we used the same growth function $g(t)=\max\{t_{0},\sqrt{t}\}$ for both LB-GP-UCB and A-GP-UCB across all experiments, where $t_{0}$ was selected so that at least 5 candidates are generated for $g(1)$. We study the impact of this choice in the ablation section. We used 10 initial points for each algorithm unless specified otherwise. To select upper bound $\theta_{0}$ for A-GP-UCB and LB-GP-UCB we fitted a length scale to initial data points with MLE (and we did not use MLE after that). We present the results in Figure 3 below. We show running times in Table 2 in Appendix I.2. We now describe each benchmark problem in detail.
@@
Berkenkamp Toy Problem We start with a one-dimensional toy problem proposed by the same paper that proposed the A-GP-UCB algorithm [6]. We showed a plot of this one-dimensional function in Figure 1. On the right side of the domain, the function appears to be smoother than on the left side. In this problem, we only use three initial points, to benchmark the ability of algorithms to escape from the local optimum on the right side of the domain. MLE and MCMC can be easily misled towards too-long length scale values, which causes them to get stuck in the local optimum. Both A-GP-UCB and LB-GP-UCB quickly find the optimal solution, however, due to over-exploration, the cumulative regret of A-GP-UCB grows faster than that of LB-GP-UCB.
@@
Michalewicz Synthetic Function As a next benchmark, we evaluate our algorithm on the five-dimensional Michalewicz synthetic function, which has been designed to be challenging while using MLE for fitting hyperparameters, because it exhibits different degrees of smoothness throughout its domain. In the histogram in Figure 2, we compare the selected length scale value with an estimate of the optimal length scale $\hat{\theta}^{*}$. We produce this estimate by sampling ten thousand points uniformly through the domain and fitting a length scale value by maximum likelihood to the points with the top 1% of objective values. In this way, we are able to capture the length scale value that produces a good model for the Michalewicz function around the optimum, where it is least smooth. We can see LB-GP-UCB selects lengthscale value closer to $\hat{\theta}^{*}$ and as a result outperforms other baselines in terms of both cumulative and best regret metrics.
@@
Material Design Problems We utilise material design tasks proposed by [17] and [30] - the 4-dimensional CrossedBarrel and 5-dimensional AGNP tasks. At each time step, the algorithm can choose which material configuration to try, and observe the objective value, which corresponds to a given material optimisation criterion. As material design problems are known to exhibit a needle-in-a-haystack behaviour [35], on both benchmarks MLE and MCMC get stuck at a suboptimal solution and their best regret does not fall beyond a certain value. A-GP-UCB is able to quickly find low-regret solutions on the AGNP benchmark, but struggles on the Crossedbarrel problem and underperforms in terms of cumulative regret. On the contrary, LB-GP-UCB performs well across both benchmark problems and across both regret metrics.
@@
Ablation on $g(t)$ To test robustness of LB-GP-UCB, we evaluate it on Michalewicz function together with A-GP-UCB for different choices of $g(t)$. We try functions of form $g(t)=\max(t_{0},t^{a})$ for $a\in(0.25,0.5,0.75)$. In Figure 4 we plot the final performance of the algorithms after $N=250$ steps as well as the distribution of selected length scale values. We see that A-GP-UCB is very sensitive to the selection of growth function $g(t)$, whereas our algorithm selects similar length scale values regardless of $g(t)$, which results in consistently good best regret results. We can also see that LB-GP-UCB typically selects values around $\hat{\theta}^{*}$ for different growth functions, whereas A-GP-UCB decreases its length scale beyond $\hat{\theta}^{*}$, resulting in slower convergence.
@@ 6 Related Work
We already mentioned the work proposing A-GP-UCB [6], which so far has been the only work providing the guarantees on BO with unknown hyperparameters, where only an upper bound on the optimal length scale value is known. The work of [39] addresses the problem, where, in addition, a lower bound on the optimal length scale is known, however, the regret bound of the algorithm they propose scales with the $\gamma_{T}(k^{\theta_{L}})$ of the smallest possible length scale $\theta_{L}$, making it no better than a naïve algorithm always selecting $\theta_{L}$. [8] studied the problem of solving BO, when the kernel function is misspecified, however, provided no method for finding the well-specified kernel function. [26] proved a lower bound on the algorithm’s regret in the case when the regularity of RKHS is unknown (which corresponds to an unknown $\nu$ hyperparameter in the case of Matérn kernel), compared to their work we focused on different unknown hyperparameters, such as the length scale. There have also been a number of works [13, 21, 27, 32] that tackled the problem of BO with unknown hyperparameters but did not provide a theoretical analysis of the used algorithm. Some of earlier works [15, 38] viewed BO with unknown hyperparameters as meta-learning or transfer learning problem, where a large dataset is available for pre-training. In our problem setting, we do not assume access to any such pre-training data. Within this work, we considered a frequentist problem setting, where the black-box function is arbitrarily selected from some RKHS. While there are no guarantees for the consistency of MLE in such a setting, if we were to assume a Bayesian setting and put a GP prior on the black box, statistical literature derived asymptotic consistency results [4, 22, 24, 28] for MLE of kernel hyperparameters, including length scale. Under such a Bayesian setting, [41] studied the problem of BO with unknown prior, when we are given a finite number of candidate priors. [10] derived predictive guarantees for GP in a Bayesian setting with unknown hyperparameters, provided that hyperpriors on those hyperparameters are known. However, the authors do not provide any BO algorithm based on their results.